# Global magnetohydrodynamic simulations of the inner regions of protoplanetary discs. I. Zero-net flux regime

### v1

@@ Abstract
The inner regions of protoplanetary discs, which encompass the putative habitable zone, are dynamically complex, featuring a well ionised, turbulent active inner region and a poorly ionised dead outer region. In this first paper, we investigate a base level model of the magnetohydrodynamic processes around this dead–active zone interface, using five three-dimensional global magnetohydrodynamic simulations in the zero-net flux regime. We employ physically motivated profiles for Ohmic resistivity and ambipolar diffusion, alongside a simplified thermodynamic model comprising a hot disc and cool corona. Our results show that, first, large-scale coherent poloidal magnetic field loops form in the magnetorotational instability (MRI) active region. These loops lead to the accumulation of tightly wound magnetic flux at the disc–corona temperature transition, driving strong, localised accretion flows in the surface layers of the active region. Second, an axisymmetric pressure maximum, that extends across multiple disc scale heights, develops as a result of outward mass transport from the active region. This triggers the Rossby wave instability and leads to the development of anticyclonic vortices, though their coherence is weakened due to the turbulent dynamics at the interface. Third, the dead zone develops magnetic field with a distinct morphology, likely resulting from the outward diffusion of large-scale poloidal loops in the active zone. This self-consistently generated field exhibits a vertical structure consistent with that required to drive accretion in the inner dead zone under the weak magnetic wind paradigm. In the second paper in the series, we extend this work to the vertical-net flux regime, where global magnetic flux transport and magnetically driven outflows become dynamically significant.
@@
pubyear: 2025pagerange: Global magnetohydrodynamic simulations of the inner regions of protoplanetary discs. I. Zero-net flux regimeGlobal magnetohydrodynamic simulations of the inner regions of protoplanetary discs. I. Zero-net flux regime
@@ 1 Introduction
The inner regions ($\lesssim 2\,\text{au}$) of protoplanetary discs encompass several key transition zones – including the star–disc boundary, silicate sublimation front, and dead–active zone interface – which significantly impact the long and short time evolution of the disc. These regions are also critically important because they include the habitable zone and are preferred sites for dust aggregation and terrestrial planet formation, regulate the protostellar assembly process through disc accretion and its associated variability, and contain the launching points of jets and magnetic outflows of different morphologies.
@@
The magnetohydrodynamic (MHD) processes operating within these regions are extremely complex. Beyond the critical radius of the dead–active zone interface ($\lesssim 1\,\mathrm{au}$), turbulence driven by the magnetorotational instability (MRI; Balbus & Hawley, 1991) is suppressed on account of the sharp drop-off in ionisation, leading to the formation of a quiescent MRI dead zone (Gammie, 1996). In this region, accretion is theorised to be driven by weak laminar magnetic pressure winds (Bai & Stone, 2013; Gressel et al., 2015). The (possible) accretion mismatch between these regions drives the formation of large-scale disc structures – rings and vortices – that can trap dust and promote planetesimal growth (e.g. Varnière & Tagger, 2006; Kretke et al., 2009). At the same time, the transition between laminar and turbulent flow imposes a mismatch in magnetic flux transport (Guilet & Ogilvie, 2014; Bai & Stone, 2017; Leung & Ogilvie, 2019). The magnetic flux influences disc evolution, by affecting the strength of MRI turbulence (Hawley et al., 1995) and the launching and efficiency of magnetic outflows (e.g. Lesur, 2021b). In addition, the mass and magnetic flux advected through the active region must ultimately be loaded onto the protostar, and interact with its dipolar magnetic field (e.g. Romanova & Owocki, 2015), further contributing to the unsteady magnetic behaviour underlying observed accretion variability (Fischer et al., 2023). Finally, the interplay between the stellar jet, episodic turbulent magnetic outflows, and laminar magnetic outflows (e.g. Tu et al., 2025), likely influences the observed morphology of chemically nested conical outflows originating from the inner disc (e.g. De Valon et al., 2020; Pascucci et al., 2024).
@@
Direct observations of the inner regions of protoplanetary discs are essential to constrain this rich panoply of physics. However, they remain extremely challenging due to the proximity of the host star, the small angular resolution required, and the high opacity at wavelengths accessible to instruments like ALMA. Nevertheless, interferometers, such as GRAVITY and MATISSE on the VLT, have begun to provide critical insights on this region. These instruments have revealed asymmetries in the inner disc of HD 163296 (e.g. Varga et al., 2021; GRAVITY Collaboration et al., 2021a), constrained the magnetospheric accretion process (GRAVITY Collaboration et al., 2024), and enabled the characterisation of hot dust near the silicate sublimation radius (GRAVITY Collaboration et al., 2021b). However, they are unable to probe the inner disc midplane directly.
@@
Therefore, despite the inherent challenges of modelling protoplanetary discs (see Haworth et al., 2016), global numerical simulations of the inner disc (e.g. Faure et al., 2014; Flock et al., 2017; Iwasaki et al., 2024, and references therein) remain indispensable for studying disc evolution and interpreting recent observational advances. Encouragingly, the upcoming generation of high-resolution radio band observation instruments – including the SKA and the ngVLA – will, by the end of the decade, provide the sensitivity and resolution needed to directly probe the inner disc and support the interpretation of theoretical models and simulations (e.g. Ilee et al., 2020; Ueda et al., 2022).
@@
This paper is the first in a series that investigates the inner regions of protoplanetary discs through exascale GPU simulations performed with the idefix code (Lesur et al., 2023b). We focus specifically on understanding the MHD processes around the dead–active zone interface using a simplified numerical setup: an initial zero-net flux (ZNF) magnetic field, physically motivated prescriptions for Ohmic and ambipolar diffusion, and a two-temperature structure representing a hot disc embedded in a cold corona. To maintain a controlled numerical setup, we omit several layers of physical complexity – such as radiative transfer (e.g. Gressel et al., 2020), dust, calculations of the ionisation fraction, and the Hall effect (e.g. Béthune et al., 2017) – that may be important in global MHD simulations. The investigation of the ZNF regime in this paper, is followed by the vertical-net flux regime in a subsequent paper, and eventually progressing toward more comprehensive models incorporating dust, the Hall effect, and the interaction with the star. This staged approach will allow us to generalise our physical insights beyond the constrained parameter space accessible to global simulations.
@@
Finally, the paper is structured as follows. In Section 2 we provide the physical and numerical background that motivates and contextualises our simulations. Section 3 outlines the inner disc model, key assumptions, and numerical methods. We then present the results of our simulations in three parts: the structure of the active zone and disc accretion in Section 4; the formation and morphology of large-scale hydrodynamic (HD) structures at the dead–active zone interface in Section 5; and the evolution of large-scale coherent magnetic fields in the dead zone in Section 6. The concluding remarks are presented in Section 7.
@@ 2 Background and inner disc physics
To establish the necessary context, we begin with a brief overview of the theory behind the dead–active zone interface in Section 2.1. Section 2.2 introduces the accretion paradigms and the role of the inner disc in driving accretion variability. This is followed by a discussion of the large-scale HD structures that emerge at the dead–active zone interface in Section 2.3. Finally, in Section 2.4, we examine the origin and global transport of large-scale magnetic flux.
@@ 2.1 The dead–active zone interface
[…]
However, a major limitation [^1] The Hall effect, which we neglect in this work, also has a significant impact on the dynamics in the dead region (e.g. Lesur et al., 2014; Béthune et al., 2017; Wurster, 2021).of this model, is that it fails to account for the increasing role of ambipolar diffusion in the low-density surface layers surrounding the dead zone, which further suppresses MRI activity (e.g. Blaes & Balbus, 1994; Kunz & Balbus, 2004). This led to a recharacterisation of the accretion paradigm in the dead zone – from one driven by MRI turbulence in the surface layers (e.g. Fleming & Stone, 2003; Turner & Sano, 2008), to one governed by weak laminar magnetic winds (e.g. Bai & Stone, 2013; Gressel et al., 2015) – as discussed further in Section 2.2.
Despite its importance, there exists a limited amount of numerical work investigating the dead–active zone interface (see the taxonomy in Section 3.2.2). Its location, shape and stability are highly complex due to the interpenetration of dynamics and thermodynamics (e.g. Wünsch et al., 2006; Latter & Balbus, 2012; Faure et al., 2014; Cecil & Flock, 2024). In the MRI active region, for instance, temperature depends on the local dissipation provided by the turbulence, whilst the turbulence depends on the temperature through the thermal ionisation. Further feedback loops also arise from dust and mass accumulation, through their influence on opacity and ionisation, all of which can drive explosive dynamics and substantial variability in the accretion rate.
@@ 2.2 Accretion and variability
An efficient and common mechanism for angular momentum is required to explain observed accretion rates of $\sim 10^{-7}$$10^{-8}\;\text{M}_{\odot}\,$yr-1 (e.g. Miotello et al., 2023) The prevailing view (see discussion in Lesur et al. 2023a) is that accretion in the poorly ionised disc ($r\gtrsim 1\,\text{au}$) is driven by the vertical extraction of angular momentum by weak laminar magnetic pressure winds (e.g. Bai & Stone, 2013; Gressel et al., 2015; Bai et al., 2016), which remain effective even at very low field strengths, such as poloidal plasma-$\beta$ values as high as $10^{8}$ (Lesur, 2021b). In contrast, accretion in the better ionised region ($r\lesssim 1\,\text{au}$) is attributed to a combination of radial turbulent stresses from MRI turbulence (e.g. Hawley et al., 1995), and radial laminar stresses from large-scale coherent magnetic fields (e.g. Zhu & Stone, 2018; Jacquemin-Ide et al., 2021). However, this paradigm critically depends on the presence of a weak poloidal magnetic field threading the disc (see Section 2.4).
@@
Protoplanetary discs also exhibit significant accretion variability in the form of strong outbursts that last for hundreds of years (e.g. FU Ori; Hartmann & Kenyon 1996). These major events obliterate the magnetospheric cavity and are often attributed to the dynamics of the dead–active zone interface. Specifically, the activation of the gravitational instability in the dead zone (Zhu et al., 2010; Bourdarot et al., 2023), allied with a reactivation of the MRI, or the possible invocation of the thermal instability (Bell & Lin, 1994; Cecil & Flock, 2024).
@@
Recently, a broad spectrum of ubiquitous, smaller-amplitude variability has been identified (see Fischer et al., 2023), with timescales linked to dynamical and viscous processes in the inner disc. This variability is classified by Fischer et al. (2023) as; routine variability (1less-than-or-similar-toabsent1$\lesssim$ 1≲ 12222 mag), occurring on timescales of hours to days in the near-IR (magnetospheric) and up to years in the mid-IR (innermost disc); bursts (1similar-toabsent1$\sim$ 1∼ 12.52.52.52.5 mag), which are distinct, discrete and last from days to a year, likely linked to stochastic instabilities associated with the star–disc connection (Stauffer et al., 2016) and episodic inner-disc processes; and large-scale outbursts, like FU Ori and EX Lup, (2.5similar-toabsent2.5$\sim$ 2.5∼ 2.56666 mag), which develop over months to years and persist for decades. Recent observations of two outburst systems, Gaia 17bpi (Hillenbrand et al., 2018) and Gaia 18dvy (Szegedi-Elek et al., 2020), provide particularly compelling evidence for the involvement of the inner disc (Cleaver et al., 2023): in both cases, the outbursts began in the near-IR before appearing in the optical, indicating an outside-in propagation originating in the inner disc.
@@
The broad range of accretion variability suggests that protoplanetary discs routinely undergo dynamical disruptions, further challenging the assumption of steady-state evolution (e.g. Terquem, 2008; Mohanty et al., 2018; Jankovic et al., 2021a, b; Delage et al., 2022). This underscores the need for global non-ideal MHD simulations of the inner disc, with careful consideration of the relevant timescales.
@@ 2.3 Large-scale hydrodynamic structure formation
In the absence of magnetic disc winds, the accretion mismatch across the dead–active zone interface inevitably leads to the formation of a localised pressure bump (e.g. Dzyurkevich et al., 2010). This pressure maximum coincides with a vortensity extremum, usually sufficient to trigger the Rossby wave instability (RWI) (Lovelace et al., 1999; Lovelace & Romanova, 2014), which saturates into coherent anticyclonic vortices. This behaviour has been confirmed in global simulations of the dead–active zone interface, including two-dimensional HD models (Lyra et al., 2009), three-dimensional unstratified MHD models (Lyra & Mac Low, 2012; Faure et al., 2014), and more advanced three-dimensional stratified MHD models incorporating realistic radiation physics (Flock et al., 2017).
These coherent structures make the dead–active zone interface a prime location for planetesimal formation, provided their formation mechanisms are robust enough to withstand the spectrum of variability intrinsic in the inner disc. For example, the accumulation of dust at the pressure maximum (e.g. Kretke et al., 2009; Ueda et al., 2017) has been proposed as a mechanism for forming multiple inner-region planets (Chatterjee & Tan, 2014; Ueda et al., 2021). Anticyclonic vortices offer an additional pathway through dust trapping in their cores (Barge & Sommeria, 1995; Tanga et al., 1996), a process extensively studied in the context of Rossby wave-induced vortices (e.g. Meheut et al., 2012; Miranda et al., 2017). However, this mechanism is only effective if the vortices do not develop internal meridional flows (Meheut et al., 2010) and are sufficiently long-lived – avoiding entering vortex life cycles (Faure et al., 2015; Flock et al., 2015) or the onset of the elliptical instability (Pierrehumbert, 1986; Lesur & Papaloizou, 2009), which can induce core turbulence or destroy the vortices.

### v3

@@ Abstract
The inner regions of protoplanetary discs, which encompass the putative habitable zone, are dynamically complex, featuring a well-ionised, turbulent active inner region and a poorly ionised ‘dead’ outer region. In this first paper, we investigate a base-level model of the magnetohydrodynamic processes around the interface between these two regions, using five three-dimensional global magnetohydrodynamic simulations in the zero-net flux regime. We employ physically motivated profiles for Ohmic resistivity and ambipolar diffusion, alongside a simplified thermodynamic model comprising a cool disc and hot corona. Our results show that, first, large-scale coherent poloidal magnetic field loops form in the magnetorotational instability active region. These loops lead to the accumulation of tightly wound magnetic flux at the disc–corona temperature transition, driving strong, localised accretion flows in the surface layers of the active region. Second, an axisymmetric pressure maximum, extending across multiple disc scale heights, develops as a result of outward mass transport from the active region. This, in turn, triggers the Rossby wave instability and leads to the development of anticyclonic vortices. Third, the dead zone develops magnetic field with a distinct morphology, likely resulting from the outward diffusion of the large-scale poloidal loops in the active zone. This self-consistently generated field exhibits a vertical structure that can drive accretion in the inner dead zone via a weak magnetic-pressure wind. In the second paper in the series, we extend this work to the vertical-net flux regime, where global magnetic flux transport and magnetically driven outflows become dynamically significant.
@@
pubyear: 2025pagerange: Global magnetohydrodynamic simulations of the inner regions of protoplanetary discs. I. Zero-net flux regimeData availability
@@ 1 Introduction
The inner regions ($\lesssim 2\,\text{au}$) of protoplanetary discs encompass several key transition zones – including the star–disc boundary, silicate sublimation front, and dead–active zone interface – that significantly impact the long and short time evolution of the disc. These regions are also critically important because they (a) include the habitable zone and are preferred sites for dust aggregation and terrestrial planet formation, (b) regulate protostellar assembly through (variable) disc accretion, and (c) contain the launching points of jets and magnetic outflows of different morphologies.
@@
The magnetohydrodynamic (MHD) processes operating within these regions are extremely complex. Beyond the critical radius of the dead–active zone interface ($\lesssim\!1\,\mathrm{au}$), turbulence driven by the magnetorotational instability (MRI; balbus_powerful_1991) is suppressed on account of the sharp drop-off in ionisation, leading to the formation of a quiescent MRI-dead zone (gammie_layered_1996). In this region, accretion is theorised to be driven by weak laminar magnetic-pressure winds (bai_wind-driven_2013; gressel_global_2015). The (possible) accretion mismatch between these regions drives the formation of large-scale disc structures – rings and vortices – that can trap dust and promote planetesimal growth (e.g. varniere_reviving_2006; kretke_assembling_2009). At the same time, the transition between laminar and turbulent flow imposes a mismatch in magnetic flux transport (guilet_global_2014; bai_hall_2017; leung_local_2019). The magnetic flux influences disc evolution, by controlling the strength of MRI turbulence (hawley_local_1995) and the launching and efficiency of magnetic outflows (e.g. lesur_systematic_2021). In addition, the mass and magnetic flux advected through the active region must ultimately be loaded onto the protostar, and interact with its dipolar magnetic field (e.g. romanova_accretion_2015), further contributing to the unsteady magnetic behaviour underlying observed accretion variability (fischer_accretion_2023). Finally, the interplay between the stellar jet, episodic turbulent magnetic outflows, and laminar magnetic outflows (e.g. tu_highly_2025), likely influences the observed morphology of chemically nested conical outflows originating from the inner disc (e.g. de_valon_alma_2020; pascucci_nested_2024).
@@
Direct observations of the inner regions of protoplanetary discs are essential to constrain this rich panoply of physics. However, they remain extremely challenging due to the proximity of the host star, the small angular resolution required, and the high opacity at wavelengths accessible to instruments like ALMA. Nevertheless, interferometers, such as GRAVITY and MATISSE on the VLT, have begun to provide critical insights on this region. For example, these instruments have revealed asymmetries in the inner disc of HD 163296 (e.g. varga_asymmetric_2021; gravity_collaboration_gravity_2021-1), constrained the magnetospheric accretion process (gravity_collaboration_gravity_2024-1), and enabled the characterisation of hot dust near the silicate sublimation radius (gravity_collaboration_gravity_2021). However, they are unable to probe the inner disc midplane directly.
@@
Therefore, despite the inherent challenges of modelling protoplanetary discs (see haworth_grand_2016), global numerical simulations of the inner disc (e.g. faure_thermodynamics_2014; flock_3d_2017; iwasaki_dynamics_2024, and references therein) remain indispensable for studying disc evolution and interpreting recent observational advances. Encouragingly, the upcoming generation of high-resolution radio band observation instruments – including the SKA and the ngVLA – will, by the end of the decade, provide the sensitivity and resolution needed to directly probe the inner disc and support/verify theoretical models and simulations (e.g. ilee_observing_2020; ueda_probing_2022).
@@
This paper is the first in a series that investigates the inner regions of protoplanetary discs through GPU simulations performed with the idefix code (lesur_idefix_2023). We focus specifically on understanding the MHD processes around the dead–active zone interface using a simplified numerical setup: an initial zero-net flux (ZNF) magnetic field, physically motivated prescriptions for Ohmic and ambipolar diffusion, and a two-temperature structure representing a cool disc embedded in a hot corona. To maintain a controlled numerical setup, we omit several layers of physical complexity – such as radiative transfer (e.g. gressel_global_2020), dust, calculations of the ionisation fraction, and the Hall effect (e.g. bethune_global_2017) – that may be important in global MHD simulations. The investigation of the ZNF regime in this paper will be followed by a study of the vertical-net flux regime in a subsequent paper, eventually progressing towards more comprehensive models incorporating dust, the Hall effect, and the interaction with the star. This staged approach will allow us to generalise our physical insights beyond the constrained parameter space accessible to global simulations.
@@
Finally, the paper is structured as follows. In Section 2 we provide the physical and numerical background that motivates and contextualises our simulations. Section 3 outlines the inner disc model, key assumptions, and numerical methods. We then present the results of our simulations in three parts: the structure of the active zone and disc accretion in Section 4; the formation and morphology of large-scale hydrodynamic (HD) structures at the dead–active zone interface in Section 5; and the evolution of large-scale coherent magnetic fields in the dead zone in Section 6. Concluding remarks are presented in Section 7.
@@ 2 Background and inner disc physics
To establish the necessary context, we begin with a brief overview of the theory underlying the dead–active zone interface in Section 2.1. Section 2.2 introduces the accretion paradigms and the role of the inner disc in driving accretion variability. This is followed by a discussion of the large-scale HD structures that emerge at the dead–active zone interface in Section 2.3. Finally, in Section 2.4, we examine the origin and global transport of large-scale magnetic flux.
@@ 2.1 The dead–active zone interface
Protoplanetary discs are weakly ionised systems, making them a unique archetype in the context of astrophysical discs (lesur_magnetohydrodynamics_2021). The seminal work of gammie_layered_1996 introduced the ‘layered accretion’ paradigm, on account of Ohmic resistivity suppressing the MRI (e.g. jin_damping_1996). This model posits that the MRI is active only in an ‘envelope’, comprising a hot inner disc ($\lesssim 0.5$$1\,\text{au}$) sufficiently ionised either thermally (e.g. umebayashi_chapter_1988) or thermionically (desch_high-temperature_2015), and disc surface layers, sufficiently ionised by stellar irradiation or cosmic rays. This envelope surrounds an MRI-dead interior. The complex interplay between the MRI, angular momentum transport, and localised accretion suggests that steady-state accretion onto the central star is unlikely within this framework (gammie_layered_1996; terquem_new_2008), implying that temporal variability is an intrinsic feature of the inner disc.
However, a major limitation [^1] The Hall effect, which we neglect in this work, also has impacts on dead-zone dynamics (e.g. lesur_thanatology_2014; bethune_global_2017; wurster_we_2021).of this model, is that it fails to account for the increasing role of ambipolar diffusion in the low-density surface layers surrounding the dead zone, which further suppresses MRI activity (e.g. blaes_local_1994; kunz_ambipolar_2004). This led to a recharacterisation of the accretion paradigm in the dead zone – from one driven by MRI turbulence in the surface layers (e.g. fleming_local_2003; turner_dead_2008), to one governed by weak laminar magnetic winds (e.g. bai_wind-driven_2013; gressel_global_2015) – as discussed further in Section 2.2.
Despite its importance, there exists a limited amount of numerical work investigating the dead–active zone interface (see the taxonomy in Section 3.2.2). Its location, shape and stability are highly complex due to the interpenetration of dynamics and thermodynamics (e.g. wunsch_two-dimensional_2006; latter_dynamics_2012; faure_thermodynamics_2014; cecil_variability_2024). For instance, in the MRI-active region, temperature depends on the local dissipation provided by the turbulence, whilst the turbulence depends on the temperature through the thermal ionisation. Further feedback loops also arise from dust and mass accumulation, through their influence on opacity and ionisation, all of which can drive explosive dynamics and substantial variability in the accretion rate.
@@ 2.2 Accretion and variability
An efficient and common mechanism for angular momentum is required to explain observed accretion rates of $\sim\!10^{-7}$$10^{-8}\;\text{M}_{\odot}\,$yr-1 (e.g. miotello_setting_2023) The prevailing view (see discussion in lesur_hydro-_2023) is that accretion in the poorly ionised disc ($r\gtrsim 1\,\text{au}$) is driven by the vertical extraction of angular momentum by weak laminar magnetic-pressure winds (e.g. bai_wind-driven_2013; gressel_global_2015; bai_magneto-thermal_2016), which remain effective even at very low field strengths, such as midplane poloidal plasma-$\beta$ values as high as $10^{8}$ (lesur_systematic_2021). In contrast, accretion in the better ionised region ($r\lesssim 1\,\text{au}$) is facilitated by a combination of radial turbulent stresses from MRI turbulence (e.g. hawley_local_1995), and radial laminar stresses from large-scale coherent magnetic fields (e.g. zhu_global_2018; jacquemin-ide_magnetic_2021). However, this paradigm critically depends on the presence of a weak poloidal magnetic field threading the disc (see Section 2.4).
@@
Protoplanetary discs also exhibit significant accretion variability in the form of strong outbursts that last for hundreds of years (e.g. FU Ori; hartmann_fu_1996). These major events obliterate the magnetospheric cavity and are often attributed to the dynamics of the dead–active zone interface, specifically, the activation of gravitational instability in the dead zone (zhu_long-term_2010; bourdarot_fu_2023), allied with a reactivation of the MRI, or possibly thermal instability (bell_using_1994; cecil_variability_2024).
@@
Recently, a broad spectrum of ubiquitous, smaller-amplitude variability has been identified (see fischer_accretion_2023), with timescales linked to dynamical and viscous processes in the inner disc. This variability is classified by fischer_accretion_2023 as: (a) routine variability ($\lesssim 1$22 mag), occurring on timescales of hours to days in the near-IR (magnetospheric) and up to years in the mid-IR (innermost disc); (b) bursts ($\sim 1$$2.5$ mag), which are distinct, discrete and last from days to a year, likely linked to stochastic instabilities associated with the star–disc connection (stauffer_csi_2016); (c) episodic inner-disc processes; and (d) large-scale outbursts, like FU Ori and EX Lup, ($\sim 2.5$66 mag), which develop over months to years and persist for decades. Recent observations of two outbursting systems, Gaia 17bpi (hillenbrand_gaia_2018) and Gaia 18dvy (szegedi-elek_gaia_2020), provide particularly compelling evidence for the involvement of the inner disc (cleaver_magnetically-activated_2023). In both cases, the outbursts began in the near-IR before appearing in the optical, indicating an outside-in propagation through the inner disc.
@@
The broad range of accretion variability suggests that protoplanetary discs routinely undergo dynamical disruptions, further challenging the assumption of steady-state evolution (e.g. terquem_new_2008; mohanty_inside-out_2018; jankovic_mri-active_2021; jankovic_mri-active_2021-1; delage_steady-state_2022). This underscores the need for global non-ideal MHD simulations of the inner disc, with careful consideration of the relevant timescales.
@@ 2.3 Large-scale hydrodynamic structure formation
In the absence of magnetic disc winds, the accretion mismatch across the dead–active zone interface inevitably leads to the formation of a localised pressure bump (e.g. dzyurkevich_trapping_2010). This pressure maximum coincides with a vortensity extremum, usually sufficient to trigger the Rossby wave instability (RWI) (lovelace_rossby_1999; lovelace_rossby_2014), which saturates in coherent anticyclonic vortices. This behaviour has been confirmed by global simulations of the dead–active zone interface, including two-dimensional HD models (lyra_planet_2009), three-dimensional unstratified MHD models (lyra_rossby_2012; faure_thermodynamics_2014), and more advanced three-dimensional stratified MHD models incorporating realistic radiation physics (flock_3d_2017).
These coherent structures make the dead–active zone interface a prime location for planetesimal formation, provided their formation mechanisms are robust enough to withstand the spectrum of variability intrinsic in the inner disc. For example, the accumulation of dust at the pressure maximum (e.g. kretke_assembling_2009; ueda_analytic_2017) has been proposed as a mechanism for forming multiple inner-region planets (chatterjee_inside-out_2014; ueda_early_2021). Anticyclonic vortices offer an additional pathway because they can trap dust in their cores (barge_did_1995; tanga_forming_1996), a process extensively studied in the context of Rossby-wave-induced vortices (e.g. meheut_dust-trapping_2012; miranda_long-lived_2017). However, this mechanism is only effective if the vortices do not develop internal meridional flows (meheut_rossby_2010) and are sufficiently long-lived – avoiding cycles of formation and destruction (faure_vortex_2015; flock_gaps_2015) or the onset of the elliptical instability (pierrehumbert_universal_1986; lesur_stability_2009), which can induce core turbulence or destroy the vortices.